# Nonparametric Proportional Hazards Model with Differential Regularization Applied to Spatial Survival Data

### v4

@@ Abstract
The Proportional Hazards (PH) model is one of the most common model used in survival analysis, which typically assumes a log-linear relationship between covariates and the hazard function. However, this assumption may not hold in practice. This paper introduces a nonparametric extension of the PH model, which generalizes the log-linear assumption by allowing for an unspecified, smooth function of covariates, enabling more flexible modeling. We focus on applications with spatial survival data, where the location of an event affects the risk. The proposed model captures this spatial variation using a nonparametric spatial effect. We estimate the spatial effect using finite element methods on a mesh constructed from a triangulation of the domain, which allows us to handle irregular shapes. The model remains within the classical partial likelihood framework, ensuring computational feasibility. To enforce the smoothness in the nonparametric spatial effect, we consider a differential penalization. We establish the asymptotic properties of the proposed estimator using sieve methods, demonstrating its consistency and the asymptotic normality of the parametric component. A simulation study is conducted to evaluate the model’s performance, followed by two empirical applications that demonstrate its practical advantages over standard PH models, especially in settings with spatial dependence in survival data.
@@ 1 Introduction
Since its introduction by (cox1972regression), the proportional hazards (PH) model has become a fundamental tool in survival analysis. Central to this framework is the hazard function, which quantifies the instantaneous risk of an event occurring at a particular time, given that the subject has not yet experienced the event up to that point.
Within the PH framework, the hazard is expressed as the product of two parts: a baseline hazard, which is the same for all individuals, and a term that incorporates the effects of covariates. The standard version of the model assumes that the covariates have a log-linear effect, meaning the logarithm of the hazard is a linear combination of the covariates. This setup provides a convenient interpretation of the regression coefficients as hazard ratios: each coefficient reflects the multiplicative change in risk associated with a one-unit increase in the corresponding covariate. These coefficients can be estimated efficiently using partial likelihood.
The assumption of log-linearity, however, can be too restrictive in practice. A more flexible alternative is the nonparametric proportional hazards model, which allows some covariate effects to be modeled in a nonparametric way rather than forcing a purely linear form. In this formulation, the hazard depends on two sets of covariates: one that enters the model linearly, with a finite set of regression coefficients, and another that enters through an unspecified smooth function. The baseline hazard is also left unspecified, which provides additional flexibility while retaining the proportional hazards structure.
Estimation approaches for the PH model that incorporate nonparametric covariate effects through smooth functions fall into two main categories. The first class consists of local likelihood methods (tibshirani1987local; fan1997local; chen2010global), which rely on kernel smoothing and the careful selection of bandwidth parameters. The second widely used class consists of spline-based approaches (o1988nonparametric; hastie1990exploring; kooperberg1995hazard), in which the smooth function is approximated using a basis expansion. This formulation enables flexible modeling while preserving parsimony and interpretability.
@@
For a comprehensive review of spatial PH models, we refer the reader to hanson2014spatial, and we restrict our discussion to the main contributions. In li2006semiparametric, the authors propose a semiparametric normal transformation model for spatial survival data. In this framework, observations marginally follow a PH model, while their joint distribution is defined by transforming the data into approximately normal variables and assuming a multivariate normal distribution for the transformed outcomes. A key limitation of this approach is that, if the transformation is misspecified, the assumption of multivariate normality may be violated, potentially leading to biased inference. Notably, this model was originally motivated by the claim that “direct nonparametric maximum likelihood estimation in such models is practically prohibited due to the high-dimensional, intractable integration in the likelihood function and the infinite-dimensional nuisance baseline hazard parameter.” This highlights the importance of basing the estimation on the partial likelihood, which, as in the standard PH model, treats the infinite-dimensional baseline hazard as a nuisance parameter, rather than attempting estimation from the full likelihood. Following this principle, the proposed model assumes independence among observations, incorporates a nonparametric spatial effect into the hazard function to capture spatial effects, and employs the partial likelihood for estimation, thereby showing that a nonparametric estimation can be carried out in a straightforward manner.
Another spatial PH alternative is the composite likelihood approach of paik2012composite, which assumes the Farlie–Gumbel–Morgenstern distribution and models the dependence parameter as a function of geographic and demographic pairwise distances. Apart from the restrictive dependence structure, composite likelihood methods—while computationally convenient—can be less efficient than full likelihood approaches, as they ignore higher-order dependencies.
Bayesian formulations have also been proposed, such as banerjee2003frailty; taylor2017, which partition the domain into clusters and assign random effects. In hennerfeind2003geoadditive, instead, a nonparametric spatial effect is modeled using splines, where spatial dependence is incorporated through the choice of priors on the spline parameters. These methods, however, rely on a parametric specification of the baseline hazard, thereby overlooking a key advantage of leaving the baseline hazard as a nuisance parameter.
@@
The smoothness of the spatial effect is enforced through a differential penalization term, which yields a concave maximization problem for a high-dimensional, differentiable objective function. This formulation enables efficient estimation using derivative-based optimization methods, facilitated by the fact that the derivative is available in closed form and simple to evaluate. A related penalization strategy for standard regression is discussed in sangalli2013spatial, although it cannot be readily extended to survival data. Another distinction is that our estimation procedure is developed within the framework of sieve theory, which enables the derivation of robust asymptotic properties.
@@ 2 Model
Let $T$ be a nonnegative time-to-event random variable representing the occurrence of the event of interest. We consider a bounded spatial domain $\Omega\subset\mathbb{R}^{2}$ with a regular boundary $\partial\Omega\in C^{2}(\mathbb{R}^{2})$. Let $\bm{X}\in\mathbb{R}^{b}$ denote a $b$-dimensional vector of covariates, and let $\bm{P}$ be a vector taking values in $\Omega$, representing the spatial location at which an observation is made. We assume that, conditional on the values $\bm{X}=\bm{x}$ and $\bm{P}=\bm{p}$, the hazard function of $T$ follows the PH model
@@
where $\lambda_{0}(t)$ is an unspecified baseline hazard function common to all observations, $\bm{\beta}_{0}\in\mathbb{R}^{b}$ is a vector of regression coefficients describing the log-linear effect of the covariates $\bm{X}$ on the hazard, and $h_{0}:\Omega\to\mathbb{R}$ is a smooth, unknown spatial effect capturing location-specific variation in risk not explained by $\bm{X}$.
@@ 3 Estimation via Penalized Likelihood and Sieve Method
To characterize the regression structure and capture spatial heterogeneity, we develop an estimation procedure based on penalized likelihood. Specifically, we assume that the observed data arise from an independent and identically distributed sample of size $n$, namely $\{(Y_{i},\delta_{i},\bm{x}_{i},\mathbf{p}_{i})\}_{i=1}^{n}.$ Our objective is to estimate the regression coefficients $\bm{\beta}_{0}$ and the spatial effect $h_{0}$ by maximizing, a penalized log-partial-likelihood functional $Q_{n}$ defined on a suitable space $\mathcal{B}\times\mathcal{H}$ as
@@
where $H^{2}_{\mathbf{n}}(\Omega):=\left\{h\in H^{2}(\Omega)\;\middle|\;\nabla h\cdot\mathbf{n}=0\text{ on }\partial\Omega\right\}$, and $M_{\mathcal{H}}$ is a constant large enough. The bound $\|h\|_{H^{2}(\Omega)}<M_{\mathcal{H}}$ is a standard regularity condition.
Maximizing the functional in (3) is an infinite-dimensional optimization problem. To make it computationally tractable, we replace the infinite-dimensional space $\mathcal{H}$ with a sequence of finite-dimensional subspaces $\mathcal{H}_{n}$ that become dense in $\mathcal{H}$ as $n\to\infty$. This is the core idea of the method of sieves (chen2007large). We use the finite element method to construct these subspaces.
@@ 3.1 The Finite Element Method
The FEM is a versatile numerical technique for approximating differential operators. Its core idea is to construct a finite-dimensional approximation space, typically spanned by locally supported basis functions. These basis functions are defined with respect to a subdivision of the computational domain $\Omega$ into simpler subdomains (quarteroni2009numerical).
@@
While the $C^{0}$ FEM construction presented above is helpful for intuition, our penalization involves the Laplacian-squared term $\int_{\Omega}(\Delta h)^{2}$, which is only well-defined for $h\in H^{2}(\Omega)$. Therefore, for estimation we employ an $H^{2}$-conforming, $C^{1}$ finite element space (e.g., the Argyris element, see brenner2008mathematical). For each triangulation $\mathcal{T}_{\eta}$ we define
$V_{\eta}:=\Big\{\,v\in C^{1}(\overline{\Omega})\;\Big|\;v|_{T}\in\mathbb{P}_{5}(T)\ \text{for every }T\in\mathcal{T}_{\eta}\,\Big\},$
@@ 3.2 Sieve Space Construction
where $B_{T}$ is the largest ball contained in triangle $T$. This condition prevents triangles from becoming arbitrarily thin, ensuring good approximation properties.
The discrete space corresponding to $H^{2}_{\mathbf{n}}(\Omega)$ is
$H_{n}:=V_{\eta(n)}\cap H^{2}_{\mathbf{n}}(\Omega).$
@@
$\mathcal{H}_{n}:=\Big\{\,h\in H_{n}\;\Big|\;\int_{\Omega}h\,d\bm{p}=0,\ \|h\|_{H^{2}(\Omega)}<M_{\mathcal{H}}\,\Big\}.$
@@
$\mathcal{J}^{\,n}h:=\mathcal{I}^{\,n}h-\frac{1}{|\Omega|}\int_{\Omega}\mathcal{I}^{\,n}h\,d\bm{p},$
where $\mathcal{I}^{\,n}$ is a $C^{1}$-conforming interpolation operator (e.g., the Argyris interpolant). Then (see, e.g., brenner2008mathematical),
@@
$\hat{\theta}_{n}=\operatornamewithlimits{argmax}_{\theta\in\Theta_{n}}Q_{n}(\theta).$ (5)
@@ 3.3 Numerical Implementation
Let $\bm{\psi}_{n}(\cdot)=(\psi_{1}(\cdot),\dots,\psi_{K(n)}(\cdot))^{\top}$ be the vector of $K(n)$ nodal basis functions for the conforming $C^{2}$ finite element space $V_{n}$. Any $h\in\mathcal{H}_{n}$ can be written as
@@
The zero-mean constraint $\int_{\Omega}h\,d\bm{p}=0$ can therefore be written as $\mathbf{1}^{\top}\mathbf{R}_{0,n}\mathbf{c}=0$. Let $\mathbf{r}_{0}=\mathbf{R}_{0,n}\mathbf{1}$ and let $\mathbf{Z}_{n}\in\mathbb{R}^{K_{n}\times(K_{n}-1)}$ be an orthonormal basis of $\ker(\mathbf{r}_{0}^{\top})$; then any feasible $\mathbf{c}$ can be written as $\mathbf{c}=\mathbf{Z}_{n}\mathbf{h}$, with $\mathbf{h}\in\mathbb{R}^{K_{n}-1}$.
@@
The objective is differentiable and concave in $(\bm{\beta},\mathbf{h})$, so derivative-based maximization applies. In fact we propose to use BFGS quasi-Newton algorithm (nocedal2006numerical) with analytic gradients. For that, define
@@
where $(\hat{\bm{\beta}},\hat{\mathbf{h}})=\operatornamewithlimits{argmax}_{(\bm{\beta},\mathbf{h})}\hat{Q}(\bm{\beta},\mathbf{h})$.
@@ 3.4 Selecting the tuning parameter
For the selection of the tuning parameter $\lambda$, we adopt a data-driven cross-validation strategy, as in cygu2021pcoxtime. In this approach, the optimal value of $\lambda$ is chosen to either minimize the cross-validated partial likelihood deviance (CV-PLD).
@@
Alternatively, the CV-C-index (harrell1996multivariable) criterion aggregates the out-of-fold risk scores and computes Harrell’s concordance index for the PH model (as implemented, for example, in the survival package; see dai2019cross). In this case, the optimal $\lambda$ maximizes the cross-validated C-index.
@@ 4 Asymptotic Analysis
We now turn to the asymptotic properties of the sieve estimator $\hat{\theta}_{n}$. Our approach follows the general framework of sieve M-estimation developed in chen2007large.
Let $d(\cdot,\cdot)$ denote a metric on the parameter space $\Theta$, defined by
@@
$\varepsilon_{n}\;:=\;\max\Big\{\,\delta_{n},\ \eta(n),\,\sqrt{\lambda_{n}}\,\Big\},\qquad\delta_{n}\asymp\sqrt{\frac{K(n)\log K(n)}{n}}.$

### v5

@@ Abstract
The Proportional Hazards (PH) model is one of the most widely used models in survival analysis, typically assuming a log-linear relationship between covariates and the hazard function. However, in the context of spatial survival data, where the time-to-event variable is associated with a spatial location within a given domain, this assumption is often unrealistic in capturing spatial effects. Thus, this paper proposes modeling the location effect through a nonparametric function of spatial location. The function is approximated using finite element methods on a triangulated mesh to accommodate irregular domains. Estimation is carried out within the classical partial likelihood framework, with smoothness of the spatial effect enforced through differential penalization. Using sieve methods, we establish the consistency and asymptotic normality of the parametric component. Simulations and two empirical applications demonstrate superior performance compared to existing approaches.
@@ 1 Introduction
Since its introduction by (Cox, 1972), the proportional hazards (PH) model has become a fundamental tool in survival analysis. Central to this framework is the hazard function, which quantifies the instantaneous risk of an event occurring at a particular time, given that the subject has not yet experienced the event up to that point.
Within the PH framework, the hazard is expressed as the product of two parts: a baseline hazard, which is the same for all individuals, and a term that incorporates the effects of covariates. The standard version of the model assumes that the covariates have a log-linear effect, meaning the logarithm of the hazard is a linear combination of the covariates. This setup provides a convenient interpretation of the regression coefficients as hazard ratios: each coefficient reflects the multiplicative change in risk associated with a one-unit increase in the corresponding covariate. These coefficients can be estimated efficiently using partial likelihood, which focuses on the ordering of event times rather than their exact distribution. This approach has the important advantage of profiling out the baseline hazard, allowing it to remain unspecified while still yielding consistent and efficient estimates of the covariate effects. As a result, inference can be conducted on hazard ratios without making parametric assumptions about the underlying hazard function.
The assumption of log-linearity, however, can be too restrictive in practice. A more flexible alternative is the nonparametric proportional hazards model, which allows some covariate effects to be modeled in a nonparametric way rather than forcing a purely linear form. In this formulation, the hazard depends on two sets of covariates: one that enters the model linearly, with a finite set of regression coefficients, and another that enters through an unspecified smooth function.
Estimation approaches for the PH model that incorporate nonparametric covariate effects through smooth functions fall into two main categories. The first class consists of local likelihood methods (Tibshirani and Hastie, 1987; Fan et al., 1997; Chen et al., 2010), which rely on kernel smoothing and the careful selection of bandwidth parameters. The second widely used class consists of spline-based approaches (O’Sullivan, 1988; Hastie and Tibshirani, 1990; Kooperberg et al., 1995), in which the smooth function is approximated using a basis expansion. This formulation enables flexible modeling while preserving parsimony and interpretability.
@@
For a comprehensive review of spatial PH models, we refer the reader to Hanson and Zhou (2014), and we restrict our discussion to the main contributions. In Li and Lin (2006), the authors propose a semiparametric normal transformation model for spatial survival data. In this framework, observations marginally follow a PH model, while their joint distribution is defined by transforming the data into approximately normal variables and assuming a multivariate normal distribution for the transformed outcomes. A key limitation of this approach is that, if the transformation is misspecified, the assumption of multivariate normality may be violated, potentially leading to biased inference. Notably, this model was originally motivated by the claim that “direct nonparametric maximum likelihood estimation in such models is practically prohibited due to the high-dimensional, intractable integration in the likelihood function and the infinite-dimensional nuisance baseline hazard parameter.” This highlights the importance of basing the estimation on the partial likelihood, which, as in the standard PH model, treats the infinite-dimensional baseline hazard as a nuisance parameter, rather than attempting estimation from the full likelihood. Following this principle, the proposed model assumes independence among observations, incorporates a nonparametric spatial effect into the hazard function to capture spatial effects, and employs the partial likelihood for estimation, thereby showing that a nonparametric estimation can be carried out in a straightforward manner.
Another spatial PH alternative is the composite likelihood approach of Paik and Ying (2012), which assumes the Farlie–Gumbel–Morgenstern distribution and models the dependence parameter as a function of geographic and demographic pairwise distances. Apart from the restrictive dependence structure, composite likelihood methods—while computationally convenient—can be less efficient than full likelihood approaches, as they ignore higher-order dependencies.
Bayesian formulations have also been proposed, such as Banerjee et al. (2003); Taylor and Rowlingson (2017), which partition the domain into clusters and assign random effects. In Hennerfeind et al. (2003), instead, a nonparametric spatial effect is modeled using splines, where spatial dependence is incorporated through the choice of priors on the spline parameters. These methods, however, rely on a parametric specification of the baseline hazard, thereby overlooking a key advantage of leaving the baseline hazard as a nuisance parameter.
@@
The smoothness of the spatial effect is enforced through a differential penalization term, which yields a concave maximization problem for a high-dimensional, differentiable objective function. This formulation enables efficient estimation using derivative-based optimization methods, facilitated by the fact that the derivative is available in closed form and simple to evaluate. A related penalization strategy for spatial regression is discussed in Sangalli et al. (2013), although it cannot be readily extended to survival data. Another distinction is that our estimation procedure is developed within the framework of sieve theory, which enables the derivation of robust asymptotic properties. Thus, our approach could also serve as a further step toward achieving asymptotic results in spatial regression with differential regularization.
@@ 2 Model
Let $T$ be a nonnegative time-to-event random variable representing the occurrence of the event of interest. We consider a compact spatial domain with non-empty interior $\Omega\subset\mathbb{R}^{2}$ with a regular boundary $\partial\Omega\in C^{2}(\mathbb{R}^{2})$. Let $\bm{X}\in\mathbb{R}^{b}$ denote a $b$-dimensional vector of covariates, and let $\bm{P}$ be a vector taking values in $\Omega$, representing the spatial location at which an observation is made. We assume that, conditional on the values $\bm{X}=\bm{x}$ and $\bm{P}=\bm{p}$, the hazard function of $T$ follows the PH model
@@
where $\lambda_{0}(t)$ is an unspecified baseline hazard function shared across all individuals; $\bm{\beta}_{0}\in\mathbb{R}^{b}$ is a vector of regression coefficients that quantifies the log-linear effects of the covariates $\bm{X}$ on the hazard; and $h_{0}:\Omega\to\mathbb{R}$ is an unknown smooth function that captures residual spatial variation in the hazard associated with location $\bm{P}$, that is not explained by the observed covariates.
@@ 3 Estimation via Penalized Likelihood and Sieve Method
In order to characterise the regression structure and capture spatial heterogeneity, we have developed an estimation procedure based on penalised likelihood. Specifically, we assume that the observed data arise from an independent and identically distributed sample of $(Y,\delta,\bm{X},\bm{P})$ of size $n$, namely $\{(Y_{i},\delta_{i},\bm{x}_{i},\mathbf{p}_{i})\}_{i=1}^{n}.$ Our objective is to estimate the regression coefficients $\bm{\beta}_{0}$ and the spatial effect $h_{0}$ by maximizing a penalized log-partial-likelihood functional $Q_{n}$ defined on a suitable space $\mathcal{B}\times\mathcal{H}$ as
@@
where $H^{2}_{\mathbf{n}}(\Omega)=\left\{h\in H^{2}(\Omega)\;\middle|\;\nabla h\cdot\mathbf{n}=0\text{ on }\partial\Omega\right\}$, and $M_{\mathcal{H}}$ is a constant large enough. The bound $\|h\|_{H^{2}(\Omega)}<M_{\mathcal{H}}$ is a consequence of the smoothness of $h_{0}$ and the fact the $\Omega$ is compact.
Maximizing the functional in (3) is an infinite-dimensional optimization problem. To make it computationally tractable, we replace the infinite-dimensional space $\mathcal{H}$ with a sequence of finite-dimensional subspaces $\mathcal{H}_{n}$ that become dense in $\mathcal{H}$ as $n\to\infty$. This is the core idea of the method of sieves (Chen, 2007). We use the finite element method to construct these subspaces.
@@ 3.1 The Finite Element Method
The FEM is a versatile numerical technique for approximating differential operators. Its core idea is to construct a finite-dimensional approximation space, typically spanned by locally supported basis functions. These basis functions are defined with respect to a subdivision of the computational domain $\Omega$ into simpler subdomains (Quarteroni and Quarteroni, 2009).
@@
While the $C^{0}$ FEM construction presented above is helpful for intuition, our penalization involves the Laplacian-squared term $\int_{\Omega}(\Delta h)^{2}$, which is only well-defined for $h\in H^{2}(\Omega)$. Therefore, for estimation we employ an $H^{2}$-conforming, $C^{1}$ finite element space (e.g., the Argyris element, see Brenner and Scott, 2008). For each triangulation $\mathcal{T}_{\eta}$ we define
$V_{\eta}=\Big\{\,v\in C^{1}(\overline{\Omega})\;\Big|\;v|_{T}\in\mathbb{P}_{5}(T)\ \text{for every }T\in\mathcal{T}_{\eta}\,\Big\},$
@@ 3.2 Sieve Space Construction
where $B_{T}$ is the largest ball contained in triangle $T$. This condition prevents triangles from becoming arbitrarily thin, ensuring good approximation properties. The discrete space corresponding to $H^{2}_{\mathbf{n}}(\Omega)$ is
$H_{n}=V_{\eta(n)}\cap H^{2}_{\mathbf{n}}(\Omega).$
@@
$\mathcal{H}_{n}=\Big\{\,h\in H_{n}\;\Big|\;\int_{\Omega}h\,d\bm{p}=0,\ \|h\|_{H^{2}(\Omega)}<M_{\mathcal{H}}\,\Big\}.$
@@
$\mathcal{J}^{\,n}h=\mathcal{I}^{\,n}h-\frac{1}{|\Omega|}\int_{\Omega}\mathcal{I}^{\,n}h\,d\bm{p},$
where $\mathcal{I}^{\,n}$ is a $C^{1}$-conforming interpolation operator (e.g., the Argyris interpolant). Then (see, e.g., Brenner and Scott, 2008),
@@
$\hat{\theta}_{n}=\operatorname*{arg\,max}_{\theta\in\Theta_{n}}Q_{n}(\theta).$ (5)
@@ 3.3 Numerical Implementation
Let $\bm{\psi}_{n}(\cdot)=(\psi_{1}(\cdot),\dots,\psi_{K(n)}(\cdot))^{\top}$ be the vector of $K(n)$ nodal basis functions for the conforming $C^{1}$ finite element space $V_{n}$. Any $h\in\mathcal{H}_{n}$ can be written as
@@
The zero-mean constraint $\int_{\Omega}h\,d\bm{p}=0$ can therefore be written as $\mathbf{1}^{\top}\mathbf{R}_{0,n}\mathbf{c}=0$. Let $\mathbf{r}_{0}=\mathbf{R}_{0,n}^{\top}\mathbf{1}$ and let $\mathbf{Z}_{n}\in\mathbb{R}^{K_{n}\times(K_{n}-1)}$ be an orthonormal basis of $\ker(\mathbf{r}_{0}^{\top})$; then any feasible $\mathbf{c}$ can be written as $\mathbf{c}=\mathbf{Z}_{n}\mathbf{h}$, with $\mathbf{h}\in\mathbb{R}^{K_{n}-1}$.
@@
The objective is differentiable and concave in $(\bm{\beta},\mathbf{h})$, so derivative-based maximization applies. In fact we propose to use BFGS quasi-Newton algorithm (Nocedal and Wright, 2006) with analytic gradients. For that, define
@@
where $(\hat{\bm{\beta}},\hat{\mathbf{h}})=\operatorname*{arg\,max}_{(\bm{\beta},\mathbf{h})}\hat{Q}(\bm{\beta},\mathbf{h})$.
@@ 3.4 Selecting the tuning parameter
For the selection of the tuning parameter $\lambda$, we adopt a data-driven cross-validation strategy, as in Cygu et al. (2021). In this approach, the optimal value of $\lambda$ is chosen to either minimize the cross-validated partial likelihood deviance (CV-PLD).
@@
Alternatively, the CV-C-index (Harrell Jr et al., 1996) criterion aggregates the out-of-fold risk scores and computes Harrell’s concordance index for the PH model (as implemented, for example, in the survival package; see Dai and Breheny, 2019). In this case, the optimal $\lambda$ maximizes the cross-validated C-index.
@@ 4 Asymptotic Analysis
We now turn to the asymptotic properties of the sieve estimator $\hat{\theta}_{n}$. Our approach follows the general framework of sieve M-estimation developed in Chen (2007). Let $d(\cdot,\cdot)$ denote a metric on the parameter space $\Theta$, defined by
@@
$\varepsilon_{n}\;=\;\max\Big\{\,\delta_{n},\ \eta(n),\,\sqrt{\lambda_{n}}\,\Big\},\qquad\delta_{n}\asymp\sqrt{\frac{K(n)\log K(n)}{n}}.$